OpenStudy (anonymous):

HELP PLEASE! What is the x-coordinate of the x-intercept of the line that passes through points (-5, 3) and (1, -3)?

OpenStudy (texaschic101):

y = mx + b we know that m is the slope and b is the y intercept. You also know y = 0 for the point it touches the x axis, because if y is anything else, your point would not be intersecting there. Thus 0 = mx + b 0 - b = mx (0 - b)/m = x now m would be your slope from the 2 sets of points in your qeustion

OpenStudy (texaschic101):

or y = mx + b....to find the x intercept, substitute 0 in for y. The slope can be found by your numbers... slope(m) = (y2 - y1)/(x2 - x1) (-5,3) x1 = -5, y1 = 3 (1,-3) x = 1 , y = -3

OpenStudy (anonymous):

thank you!!

OpenStudy (texaschic101):

your welcome :)